Please answer part a and b of the problem
5 researcher estimated that the price elasticity of demand for automobiles in the United States is -1.2, while the income elasticity of demand is 3.0. Next year, U.S. automakers intend to increase the average price of automobiles by 5 percent, and they expect consumers disposable income to rise by 3 percent. (a) If sales of domestically produced automobiles are 8 million this year, how many automobiles do you expect U.S. automakers to sell next year? (b) By how much should domestic automakers increase the price of automobiles if they wish to increase sales by 5 percent next year?
Answer #1

Given that,

Average price increase by 5%

And elasticity of demand is -1.2

So there will be decrease in demand of 5*1.2 = 6%

Income increase by 3%

Elasticity of income is 3 So there will be increase in income 3*3 = 9%

So effectively, there will be net change in demand due to increase in income is 9-6 = 3%

Now, last year sales = 8million

So current year expected sales = 8 + 8*3% = 8.24 million

b)

If auto maker want to increase the sales, then price should be decreased.

Elasticity = % change in demand / % change in price

-1.2 = 5%/% change in price

% change in price = 5%/-1.2

% change in price = -4.17%